OpenStudy (anonymous):

\[\sqrt{12a^5b}*\sqrt{8a^6b^4}\]

OpenStudy (saifoo.khan):

\[\Huge 4\sqrt{6a^{11}b^5}\]

OpenStudy (anonymous):

not done yet are you?

OpenStudy (anonymous):

\[4a^5b^2\sqrt{6ab}\]

OpenStudy (saifoo.khan):

\begin{array}l\color{red}{\text{A}}\color{orange}{\text{g}}\color{#9c9a2e}{\text{a}}\color{green}{\text{i}}\color{blue}{\text{n}}\color{purple}{\text{ }}\color{purple}{\text{i}}\color{red}{\text{n}}\color{orange}{\text{c}}\color{#9c9a2e}{\text{o}}\color{green}{\text{r}}\color{blue}{\text{r}}\color{purple}{\text{e}}\color{purple}{\text{c}}\color{red}{\text{t}}\color{orange}{\text{?}}\color{#9c9a2e}{\text{?}}\color{green}{\text{ }}\color{blue}{\text{:}}\color{purple}{\text{s}}\color{purple}{\text{}}\end{array}

OpenStudy (saifoo.khan):

are u sure satellite?

OpenStudy (anonymous):

4a^5b^2*sqrt(6ab)

OpenStudy (anonymous):

yes i am sure. you cannot have a power higher than one inside a square root

OpenStudy (anonymous):

i mean you can, but it is not in simplest radical form.

OpenStudy (saifoo.khan):

OIC, i got u now! :)

OpenStudy (anonymous):

sqrt{5y+1}=7

OpenStudy (anonymous):

\[\sqrt{5y+1}=7\]

OpenStudy (saifoo.khan):

5y + 1 = 49

OpenStudy (anonymous):

5y+1=49 5y=48 y=9.6

OpenStudy (anonymous):

thanks!

OpenStudy (saifoo.khan):

i was correct too!

hero (hero):

lol, I saw that Saifoo

OpenStudy (anonymous):

\[(9-\sqrt{10)}(5+\sqrt{10)}\]

OpenStudy (saifoo.khan):

:S

OpenStudy (anonymous):

would this be 4@RT{10}+35

OpenStudy (saifoo.khan):

\[\Huge \color\purple {35+4\sqrt{10}}\]

OpenStudy (anonymous):

\[4\sqrt{10}+35?\]

OpenStudy (anonymous):

YESSSS!!! LOL THANK YOU!

OpenStudy (saifoo.khan):

hehe!

hero (hero):

Well Saif, at least you got something right Guru

OpenStudy (anonymous):

lol

OpenStudy (saifoo.khan):

Fight like these?? http://www.youtube.com/watch?v=jUi68aPYP5k

OpenStudy (anonymous):

\[\sqrt[3]{56?}\div \sqrt[3]{7?}\]

OpenStudy (anonymous):

would this one be 2?

OpenStudy (anonymous):

LMAO @ VIDEO LOL

OpenStudy (saifoo.khan):

Hehe.

OpenStudy (anonymous):

hahaha!!!!

OpenStudy (saifoo.khan):

2

hero (hero):

lame!

OpenStudy (saifoo.khan):

\begin{array}l\color{red}{\text{T}}\color{orange}{\text{W}}\color{#9c9a2e}{\text{O}}\color{green}{\text{!}}\color{blue}{\text{}}\end{array}

OpenStudy (anonymous):

fatman can't chop the board with his foot smh

OpenStudy (saifoo.khan):

LOL.

OpenStudy (anonymous):

he was trying so hard! lol

hero (hero):

That's because he thought his foot was stronger than his hands

OpenStudy (anonymous):

losing! lol...chopping with his head-WINNING! LOL

OpenStudy (anonymous):

9swrt y+5sqrt y